Could you please help me with this one too
(3/2) ^ x = 16/81

Answer by stanbon(48569)

(3/2) ^ x = 16/81
Rewrite as follows:
(2/3)^(-x)=(2/3)^4
-x=4
x=-4
Cheers,
Stan H.